OpenStudy (anonymous):

solve equation: 2(2)^(x-8)-2=14 express answer using log

OpenStudy (anonymous):

add two to both sides divide both sides by 2 you do not need logs for this

OpenStudy (anonymous):

2/2(2)^(x-8)=16/2 2^(x-8)=8

OpenStudy (mertsj):

Now write 8 as 2^3

OpenStudy (anonymous):

2^(x-8)=2^3

OpenStudy (mertsj):

Now. The expressions are equal, the bases are the same so the exponents must be the same. Write and solve that equation.

OpenStudy (anonymous):

x-8=3 x=11

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Thank you!